Consider an aluminium wire with a diameter of 1.07 mm. What would be the diameter of a copper wire of the same length with the same resistance? Would it be more or less massive than the aluminium wire and by what factor? The density of copper is roughly 8,960 kgm^(−3).
Answer #1

Here rho sign stands for resistivity.
Answer #2

Given:

  • Diameter of aluminium wire, dAl=1.07mm

  • Resistivity of aluminium, ρAl=2.65×108Ωm

  • Resistivity of copper, ρCu=1.68×108Ωm

  • Density of copper, ρmass, Cu=8960kg/m3

  • Density of aluminium, ρmass, Al=2700kg/m3

Step 1: Equal Resistance Condition

Resistance R for a wire is given by:

R=ρLA

For equal resistance and length:

ρAlAAl=ρCuACuACu=AAlρCuρAl

Step 2: Calculate Cross-Sectional Area

AAl=πdAl24=π(1.07×103)24=8.99×107m2ACu=8.99×1071.68×1082.65×108=5.70×107m2

Step 3: Find Copper Wire Diameter

dCu=4ACuπ=4×5.70×107π=0.853mm

Step 4: Compare Masses

Mass is given by m=ρmassAL. For the same length L:

mCumAl=ρmass, CuACuρmass, AlAAl=8960×5.70×1072700×8.99×107=2.10

Answer:

  • Diameter of copper wire: 0.853mm

  • Mass comparison: The copper wire is 2.10 times more massive than the aluminium wire


answered by: anonymous